Your firm currently has $ 80 million in debt outstanding with an 8 % interest rate. The terms of the loan require the firm to repay

$ 20 million of the balance each year. Suppose that the marginal corporate tax rate is 21%​, and that the interest tax shields have the same risk as the loan. What is the present value of the interest tax shields from this​ debt?
Answer #1

interest cost savings in year 1

=80*8%*21%

=1.3440

interest cost savings in year 2

=60*8%*21%

=1.0080

interest cost savings in year 3

=40*8%*21%

=0.6720

interest cost savings in year 4

=20*8%*21%

=0.3360

What is the present value of the interest tax shields from this​ debt

=1.3440/(1+8%)^1+1.0080/(1+8%)^2+0.6720/(1+8%)^3+0.3360/(1+8%)^4

=2.89 million

the above is answer..